Jill's age next year will be twice Ann's age last year. Their present ages total 54. How old are they now?
Answer by mananth(16946)

jills age now = x
Ann's age now = y
x+y=54................1
...
jill next year x+1
Ann last year=y-1
x+1=2(y-1)
x+1=2y-2
x-2y=-3................2
multiply 1 by 2
2x+2y=108
add to 2
3x=105
x=35 Jills age
Ann's age = 54-35=19
,,,
m.ananth@hotmail.ca